# Learning New Tasks from a Few Examples with Soft-Label Prototypes

### v3

@@ 10.1 Individual vs joint prediction
where $\mathbf{F(x)}$ is the ensemble and $\mathbf{f(x)}$ are the constituent models. The idea is that the probability distribution of the ensemble fits the target distribution more closely than the corresponding expected probability distributions of its constituent models. For joint soft-label prototypes parameterised by $g_{1}$ and $g_{2}$ and located at $p_{1}$ and $p_{2}$, we have:
@@
We plot the negative log likelihoods for Airline and Disaster on the test set after finetuning each model on ten subsets of few-shot training data (as explained before in Section 5) in Figure 6 to assess whether DeepSLP exhibits this property of ensemble methods.
We find that $NLL(\mathbf{F(x)},y)$ is much lower than $\mathbb{E}[NLL(\mathbf{f(x)},y)]$, which confirms that the joint utilisation of prototypes results in better predictions than if they were to be used individually. Our experiments confirm that this is a general trend we observe across tasks. In the following sections, we investigate the reasons behind the high values observed for $\mathbb{E}[NLL(\mathbf{f(x)},y)]$, and compare the jointly trained prototypes against a strong baseline, the fine-tuned BERT baseline.
10.2 Jointly utilised soft-label prototypes improve diversity
@@
If we use the Jenson-Shannon divergence as a diversity measure for an ensemble given by $JSD_{p(f)}[y|f(x)]=\frac{1}{M}\sum KL[y|f(x)||y|F(x)]$ where KL is the average KL divergence between the output distribution of each soft-label prototype and the jointly utilised soft-label prototypes, from Abe et al. (2022), this expression reduces to:
$H([y|F(x)]=\stackrel{{\scriptstyle\underline{ens.\;diversity}}}{{JSD_{p(f)}[y|f(x)]}}+\stackrel{{\scriptstyle\underline{avg.\;model\;uncert.}}}{{E_{p(f)}[H[y|f(x)]]}}$
@@
The above provide evidence that our approach as a whole exhibits desirable properties of ensembles which drive a higher performance but which do not lead to a higher training time nor compute.
11 Conclusion
We developed a novel few-shot learning approach for data lean settings that is based on soft-label prototypes capturing the simultaneous membership of data points over several classes. We presented two variants, one that generates fixed prototypes via solving for linear constraints, and another that learns and generates them dynamically within a neural gradient-based optimisation approach. We evaluated our methods on 48 different tasks / settings and experimentally demonstrated that our neural approach outperforms a range of strong baselines. Finally, we presented an in-depth analysis of the characteristics of our approach and analysed the reasons behind its effectiveness.

### v4

@@ A.8.1 Individual vs joint prediction
where $F(x)$ is the ensemble and $f(x)$ are the constituent models. The idea is that the probability distribution of the ensemble fits the target distribution more closely than the corresponding expected probability distributions of its constituent models. For joint soft-label prototypes parameterised by $g_{1}$ and $g_{2}$ and located at $p_{1}$ and $p_{2}$, we have:
@@
We plot the negative log likelihoods for Airline and Disaster on the test set after fine-tuning each model on ten subsets of few-shot training data (as explained before in Section 4) in Figure 8 to assess whether DeepSLP exhibits this property of ensemble methods. We find that $NLL(\mathbf{F(x)},y)$ is much lower than $\mathbb{E}[NLL(\mathbf{f(x)},y)]$, which confirms that the joint utilisation of prototypes results in better predictions than if they were to be used individually. Our experiments confirm that this is a general trend we observe across tasks. In the following sections, we investigate the reasons behind the high values observed for $\mathbb{E}[NLL(\mathbf{f(x)},y)]$, and compare the jointly trained prototypes against a strong baseline, the fine-tuned BERT baseline.
A.8.2 Jointly utilised soft-label prototypes improve diversity
@@
If we use the Jenson-Shannon divergence as a diversity measure for an ensemble given by
$JSD_{p(f)}[y|f(x)]=\frac{1}{M}\sum KL[y|f(x)||y|F(x)]$
where KL is the average KL divergence between the output distribution of each soft-label prototype and the jointly utilised soft-label prototypes, from Abe et al. (2022), this expression reduces to: ${\textstyle H([y|F(x)]=\stackrel{{\scriptstyle\underline{ens.\;diversity}}}{{JSD_{p(f)}[y|f(x)]}}+\stackrel{{\scriptstyle\underline{avg.\;model\;uncert.}}}{{E_{p(f)}[H[y|f(x)]]}}}$
@@
The above provide evidence that our approach as a whole exhibits desirable properties of ensembles which drive a higher performance but which do not lead to higher training time nor compute.